# TaP: A Taxonomy-Guided Framework for Automated and Scalable Preference Data Generation

### v3

@@ 3.4 Fine-grained Preference Annotation
We employ multiple LLMs to perform preference annotations for each pairwise combination of selected responses to a given prompt. Rather than relying on a single model, each response pair is annotated by at least five different LLLMs. Specifically, we use GPT-4 together with five high-performing open-source LLMs, including Qwen2-72B-Instruct, Command R+, Mistral-Large-Instruct-2407, Llama-3.1-70B-Instruct, and DeepSeek V2.5, to generate independent preference judgments for all pairwise comparisons. [^6] Due to the high cost of using GPT-4 for preference annotation, it is used only to annotate pairwise responses to prompts it generates. In contrast, annotations for responses to DeepSeek-V2 prompts are handled by open-source LLMs.
@@
For preference annotation, we adopt a pairwise comparison approach, wherein LLMs are prompted to choose the preferred response between two alternatives, rather than assigning independent scores to each response as in pointwise grading. Our preliminary experiments indicate that pointwise grading often fails to distinguish subtle differences between similar responses, frequently assigning them identical scores. Despite its higher annotation cost, pairwise comparison is employed to obtain more precise preference annotations. In this approach, we present two responses to LLMs and ask them to score each one across four dimensions: relevance, correctness, clarity, and completeness. These dimensions facilitate a comprehensive assessment of response quality. The resulting scores can be used to train fine-grained reward models, which in turn enhance control over LLM-generated responses by assigning weights to rewards from different fine-grained reward models during RLHF (DBLP:conf/nips/WuHSDSASOH23). The prompt used for pairwise comparison is provided in Figure 9 of Appendix C.1. Additionally, to mitigate position bias during pairwise comparison (DBLP:conf/acl/WangLCCZLCKLLS24), we adopt the Balanced Position Calibration strategy, which involves conducting an additional comparison with the positions of the two responses swapped.
To derive preference labels from the pairwise comparison scores, we compute an overall score for each response in a pair. The response with the higher overall score is labeled as preferred. To compute each overall score, we first average the scores for each evaluation dimension across both positions to obtain calibrated scores. These calibrated scores are then equally weighted and averaged to produce a final score for each response. Although domain-specific applications might prioritize certain dimensions, for example, in medical applications, correctness may be prioritized, we assign equal weight to all four dimensions to support general-purpose LLM training.
@@ 4 Dataset Statistics
We employ GPT-4 and DeepSeek-V2 to generate prompts corresponding to the 724 subjects defined in our taxonomy, which is derived from the Undergraduate Program Catalog of Regular Higher Education Institutions of China. GPT-4 generates a total of 27,715 prompts, while DeepSeek-V2 produces 50,929 prompts. The distributions of these prompts across the 12 primary disciplines in the taxonomy are presented in Figure 2. As shown in Figure 2, despite using different LLMs for prompt generation, the resulting prompt distributions are highly consistent, indicating that the TaP framework enables fine-grained control over dataset composition. In addition, since some prompts need to be regenerated after failing the feasibility check, a single question type may correspond to multiple prompts. As a result, GPT-4 yields 19,840 distinct question types, and DeepSeek-V2 yields 24,839 distinct question types.
For constructing the preference dataset, we rely solely on preference annotations generated by GPT-4, given its pioneering performance. [^7] While incorporating annotations from multiple LLMs might improve dataset quality, we leave the integration of annotations from other open-source LLMs for future work. Table 5 presents a comparative summary of the preference dataset constructed by TaP and existing open-source preference datasets. Compared to these existing datasets, the TaP-constructed dataset encompasses a larger number of samples and prompts, features longer prompts and responses, and provides fine-grained annotations.
@@ 5 Annotation Agreement
To assess the reliability of preference annotations produced by LLMs, we examine three distinct aspects of annotation agreement: (i) agreement among human annotators, (ii) agreement between human annotators and LLMs, and (iii) agreement across different LLMs. To this end, we randomly sample 300 prompts along with their corresponding paired responses generated by two distinct LLMs, as well as the associated preference annotations produced by the models. We then recruit three PhD students, all native Chinese speakers with research interests in computer science, to independently annotate the preferred response for each response pair. Label Studio [^8] https://github.com/HumanSignal/label-studio is employed to provide an annotation interface that facilitates the annotation process. For each prompt, annotators are presented with the prompt and the two corresponding LLM-generated responses (Response A and Response B) and are instructed to indicate their preference by selecting one of three options: “Response A Win,” “Response B Win,” or “Tie.”
Table 2 reports the pairwise annotation agreement across different annotators on 300 randomly selected samples. The agreement among human annotators approaches 80%, exceeding the 60%-75% range commonly reported in prior work (DBLP:journals/corr/abs-2204-05862; DBLP:conf/nips/Ouyang0JAWMZASR22; DBLP:conf/icml/CuiY0YH0NXXL0024). Furthermore, the agreement between human annotators and LLM annotators is close to or, in some cases, exceeds 80%, reaching up to 90.33%, which suggests that preference annotations generated by LLMs reliably reflect human preferences. Notably, the agreement among different LLM annotators consistently surpasses 90%, substantially outperforming the level of agreement observed among human annotators.
When computing pairwise preference annotation agreement, we observe a small number of exceptional cases in which LLMs fail to strictly follow the annotation instructions and do not produce a valid preference score, preventing the extraction of a corresponding preference label. Such samples are therefore excluded from the agreement computation. Importantly, these cases are extremely rare: among the 300 randomly selected samples, at most seven LLM-generated annotations fail to yield a valid preference label. As a result, excluding these samples has a negligible impact on the overall results and does not affect the validity of our experimental conclusions.
6 Experiments
To assess the effectiveness of TaP and to evaluate the quality of the preference datasets it constructs, we conducted both supervised fine-tuning and preference fine-tuning, including PPO and DPO, on LLMs using the datasets generated by TaP.
6.1 Supervised Fine-tuning
To conduct supervised fine-tuning using the preference data generated by TaP, we utilized the prompts from the dataset along with GPT-4-generated responses. To ensure a balanced distribution of question types, we randomly selected one prompt and its corresponding response for each question types with multiple prompts. The resulting dataset comprises 19,840 samples generated by GPT-4 and 24,839 samples produced by DeepSeek-V2. A comparative summary of this supervised fine-tuning dataset, alongside other open-source Chinese datasets, is provided in Table 3.
6.2 Preference Fine-tuning
We adopt two dominant approaches for preference fine-tuning of LLMs using our constructed preference dataset: PPO and DPO. To facilitate comparison with other open-source preference datasets, we utilize Huozi-RLHF (huozi) and Chinese-DPO-Pairs [^9] https://huggingface.co/datasets/wenbopan/Chinese-dpo-pairs for preference fine-tuning. Further details regarding the experimental setup for PPO and DPO can be found in Appendix D.1.
@@ 6.2 Preference Fine-tuning
Alpaca-GPT-4-ZH (DBLP:journals/corr/abs-2304-03277) 48,818 Chinese
COIG (DBLP:journals/corr/abs-2304-07987) 178,246 Chinese
MOSS-SFT (MIR-2023-12-294) 1,074,551 English, Chinese
COIG-CQIA (DBLP:journals/corr/abs-2403-18058) 44,694 Chinese
@@
BELLE-SFT (BELLE) 3,606,402 Chinese
@@
Table 3: Comparison of open-source datasets containing a substantial number of samples in Chinese for supervised fine-tuning with datasets derived from preference datasets constructed by TaP. “TaP-SFT (GPT-4)” and “TaP-SFT (DeepSeek-V2)” indicate that the prompts were generated by GPT-4 and DeepSeek-V2. For the Infinity-Instruct dataset, which includes samples in multiple languages, Chinese samples were selected based on the language tags provided within the data.
@@
Alpaca-GPT-4-ZH (DBLP:journals/corr/abs-2304-03277) 4.16 4.87 4.51 3.99 4.75 4.37 4.83 4.68
COIG (DBLP:journals/corr/abs-2304-07987) 2.73 4.12 3.42 2.84 4.03 3.43 3.35 3.44
MOSS-SFT (MIR-2023-12-294) 3.82 4.96 4.39 3.73 4.82 4.27 4.66 4.38
COIG-CQIA (DBLP:journals/corr/abs-2403-18058) 3.61 4.38 3.99 3.65 4.31 3.98 4.47 4.35
@@
Alpaca-GPT-4-ZH (DBLP:journals/corr/abs-2304-03277) 4.90 5.31 5.10 4.53 5.14 4.83 5.02 4.89
COIG (DBLP:journals/corr/abs-2304-07987) 2.82 4.41 3.62 3.11 4.38 3.75 3.48 3.66
MOSS-SFT (MIR-2023-12-294) 4.71 5.26 4.98 4.55 5.17 4.86 5.06 4.90
COIG-CQIA (DBLP:journals/corr/abs-2403-18058) 4.72 5.13 4.92 4.51 5.02 4.76 4.73 4.51
@@
Alpaca-GPT-4-ZH (DBLP:journals/corr/abs-2304-03277) 2.77 4.28 3.53 2.64 4.12 3.38 4.21 4.28
COIG (DBLP:journals/corr/abs-2304-07987) 1.79 2.98 2.38 1.88 3.04 2.46 2.68 2.74
MOSS-SFT (MIR-2023-12-294) 2.84 4.41 3.63 2.69 4.32 3.50 4.21 4.06
COIG-CQIA (DBLP:journals/corr/abs-2403-18058) 2.54 3.57 3.06 2.39 3.56 2.97 3.32 3.34
@@
Alpaca-GPT-4-ZH (DBLP:journals/corr/abs-2304-03277) 3.17 4.22 3.70 3.19 4.17 3.68 4.39 4.09
COIG (DBLP:journals/corr/abs-2304-07987) 1.70 2.81 2.25 1.95 2.80 2.38 2.69 2.65
MOSS-SFT (MIR-2023-12-294) 2.84 4.21 3.53 2.73 4.19 3.46 4.14 4.23
COIG-CQIA (DBLP:journals/corr/abs-2403-18058) 2.69 3.58 3.14 2.64 3.56 3.10 3.36 3.13
@@
Alpaca-GPT-4-ZH (DBLP:journals/corr/abs-2304-03277) 5.64 5.73 5.68 5.34 5.73 5.54 5.44 5.36
COIG (DBLP:journals/corr/abs-2304-07987) 3.30 4.84 4.07 3.38 4.86 4.12 3.63 3.56
MOSS-SFT (MIR-2023-12-294) 5.25 5.59 5.42 4.96 5.48 5.22 5.26 5.04
COIG-CQIA (DBLP:journals/corr/abs-2403-18058) 5.14 5.27 5.20 4.94 5.26 5.10 4.92 4.78
@@
Table 4: Performance comparison of five LLMs on AlignBench and MT-Bench-zh after supervised fine-tuning with various datasets. “GPT-4o” and “DeepSeek-V3” indicate that the responses for the respective benchmarks were evaluated using GPT-4o and DeepSeek-V3, respectively.
@@
Qwen2.5-3B-SFT-Open Huozi-RLHF (huozi) 4.02 5.23 4.63 3.88 5.04 4.46 5.03 4.94
@@
Qwen2.5-3B-SFT-TaP Huozi-RLHF (huozi) 5.15 5.82 5.49 4.65 5.50 5.08 5.72 5.24
@@
Qwen2.5-7B-SFT-Open Huozi-RLHF (huozi) 4.53 5.39 4.96 4.37 5.34 4.85 5.25 4.98
@@
Qwen2.5-7B-SFT-TaP Huozi-RLHF (huozi) 6.24 6.34 6.29 5.83 6.00 5.92 6.10 5.85
@@
Llama-3.1-8B-SFT-Open Huozi-RLHF (huozi) 3.52 5.19 4.36 3.29 5.06 4.17 4.85 4.58
@@
Llama-3.1-8B-SFT-TaP Huozi-RLHF (huozi) 3.70 5.22 4.46 3.33 4.86 4.09 5.19 4.84
@@
Gemma-2-9B-SFT-Open Huozi-RLHF (huozi) 3.93 5.22 4.58 3.54 5.05 4.29 4.77 4.62
@@
Gemma-2-9B-SFT-TaP Huozi-RLHF (huozi) 4.18 5.23 4.70 3.77 4.93 4.35 5.39 5.01
@@
Qwen2.5-14B-SFT-Open Huozi-RLHF (huozi) 5.75 5.90 5.82 5.64 5.81 5.73 5.41 5.48
@@
Qwen2.5-14B-SFT-TaP Huozi-RLHF (huozi) 6.94 6.64 6.79 6.41 6.28 6.35 6.25 6.11
@@
Qwen2.5-3B-SFT-Open Huozi-RLHF (huozi) 4.13 4.88 4.51 4.44 5.36 4.90 5.25 5.20
@@
Qwen2.5-3B-SFT-TaP Huozi-RLHF (huozi) 4.61 5.27 4.94 4.62 5.55 5.08 5.86 5.76
@@
Qwen2.5-7B-SFT-Open Huozi-RLHF (huozi) 4.71 5.56 5.13 4.46 5.43 4.95 5.21 5.19
@@
Qwen2.5-7B-SFT-TaP Huozi-RLHF (huozi) 6.24 6.44 6.34 5.67 5.98 5.82 6.25 6.19
@@
Llama-3.1-8B-SFT-Open Huozi-RLHF (huozi) 3.72 5.27 4.50 3.41 4.97 4.19 4.93 4.89
@@
Llama-3.1-8B-SFT-TaP Huozi-RLHF (huozi) 3.94 5.25 4.60 3.61 4.86 4.24 5.25 5.04
@@
Gemma-2-9B-SFT-Open Huozi-RLHF (huozi) 3.75 4.75 4.25 3.81 5.14 4.47 4.87 4.85
@@
Gemma-2-9B-SFT-TaP Huozi-RLHF (huozi) 3.85 5.20 4.53 3.56 4.96 4.26 5.27 5.11
@@
Qwen2.5-14B-SFT-Open Huozi-RLHF (huozi) 5.61 5.90 5.75 5.69 5.93 5.81 5.78 5.78
@@
Qwen2.5-14B-SFT-TaP Huozi-RLHF (huozi) 6.61 6.50 6.55 6.42 6.34 6.38 6.41 6.51

### v4

@@ 3.4 Fine-grained Preference Annotation
We employ multiple LLMs to perform preference annotations for each pairwise combination of selected responses to a given prompt. Rather than relying on a single model, each response pair is annotated by at least five different LLMs. Specifically, we use GPT-4 together with five high-performing open-source LLMs, including Qwen2-72B-Instruct, Command R+, Mistral-Large-Instruct-2407, Llama-3.1-70B-Instruct, and DeepSeek V2.5, to generate independent preference judgments for all pairwise comparisons. [^6] Due to the high cost of using GPT-4 for preference annotation, we use it only to annotate response pairs for prompts it generates. In contrast, response pairs for prompts generated by DeepSeek-V2 are annotated by open-source LLMs.
@@
For preference annotation, we adopt a pairwise comparison approach, wherein LLMs choose the preferred response between two alternatives, rather than assigning independent scores to each response as in pointwise grading. Our preliminary experiments indicate that pointwise grading often fails to distinguish subtle differences between similar responses, frequently assigning them identical scores. Despite its higher annotation cost, we employ pairwise comparison to obtain more precise preference annotations. In this approach, we present two responses to LLMs and ask them to score each one across four dimensions: relevance, correctness, clarity, and completeness. These dimensions facilitate a comprehensive assessment of response quality. The resulting scores can be used to train fine-grained reward models, which in turn enhance control over LLM-generated responses by assigning weights to rewards from different fine-grained reward models during RLHF (Wu et al., 2023). The prompt used for pairwise comparison is provided in Figure 9 of Appendix C.1. Additionally, to mitigate position bias in pairwise comparison, where LLMs tend to assign higher scores to the response appearing in a particular position (e.g., the first or second slot), we adopt the Balanced Position Calibration strategy (Wang et al., 2024a), which involves conducting an additional comparison with the positions of the two responses swapped.
To derive preference labels from pairwise comparison scores, we compute an overall score for each response and label the higher-scoring response as preferred. Specifically, we average the score for each evaluation dimension across both positions (original vs. swapped) to obtain calibrated scores. We then assign equal weights to the calibrated dimension scores and average them to produce a final score for each response. Although domain-specific applications may prioritize certain dimensions (e.g., correctness in medical applications), we use uniform weights across all four dimensions to support general-purpose LLM training.
@@ 4 Dataset Statistics
We employ GPT-4 and DeepSeek-V2 to generate prompts for the 724 subjects in our taxonomy, derived from the Undergraduate Program Catalog of Regular Higher Education Institutions of China. GPT-4 generates 27,715 prompts, whereas DeepSeek-V2 produces 50,929 prompts. The distributions of these prompts across the 12 primary disciplines are presented in Figure 2. As shown in Figure 2, despite using different LLMs for prompt generation, the resulting distributions are highly consistent, indicating that TaP enables fine-grained control over dataset composition. In addition, because some prompts must be regenerated after failing the feasibility check, a single question type may correspond to multiple prompts. Consequently, GPT-4 yields 19,840 distinct question types, and DeepSeek-V2 yields 24,839 distinct question types.
For constructing the preference dataset, we rely solely on preference annotations generated by GPT-4, given its pioneering performance. [^7] While incorporating annotations from multiple LLMs might improve dataset quality, we leave the integration of annotations from other open-source LLMs for future work. Table 5 presents a comparative summary of the preference dataset constructed by TaP and existing open-source preference datasets. Compared with these datasets, the TaP-constructed dataset contains more samples and prompts, includes longer prompts and responses, and provides fine-grained annotations.
@@ 5 Annotation Agreement
To assess the reliability of preference annotations produced by LLMs, we examine three aspects of annotation agreement: (i) agreement among human annotators, (ii) agreement between human annotators and LLMs, and (iii) agreement across different LLMs. To this end, we randomly sample 300 prompts, their paired responses generated by two distinct LLMs, and the corresponding model-generated preference annotations. We then recruit three PhD students, all native Chinese speakers with research interests in computer science, to independently label the preferred response for each response pair. Label Studio [^8] https://github.com/HumanSignal/label-studio is employed to provide an annotation interface that facilitates the process. For each prompt, annotators are shown the prompt and the two LLM-generated responses (Response A and Response B) and are instructed to indicate their preference by selecting one of three options: “Response A Win,” “Response B Win,” or “Tie.”
@@
Alpaca-GPT-4-ZH (Peng et al., 2023) 48,818 Chinese
COIG (Zhang et al., 2023a) 178,246 Chinese
MOSS-SFT (Sun et al., 2024) 1,074,551 English, Chinese
COIG-CQIA (Bai et al., 2024) 44,694 Chinese
@@
BELLE-SFT (BELLEGroup, 2023) 3,606,402 Chinese
@@
Table 3: Comparison of open-source datasets containing a substantial number of Chinese samples for supervised fine-tuning with datasets derived from TaP-constructed preference datasets. “TaP-SFT (GPT-4)” and “TaP-SFT (DeepSeek-V2)” indicate that the prompts were generated by GPT-4 and DeepSeek-V2, respectively. For the multilingual Infinity-Instruct dataset, Chinese samples were selected based on language tags in the data.
Table 2 reports pairwise annotation agreement among different annotators on 300 randomly selected samples. Agreement among human annotators approaches 80%, exceeding the 60%–75% range commonly reported in prior work (Bai et al., 2022; Ouyang et al., 2022; Cui et al., 2024). Furthermore, agreement between human annotators and LLM annotators is close to, and in some cases exceeds, 80%, reaching 90.33%. This suggests that preference annotations generated by LLMs reliably reflect human preferences. Notably, agreement among LLM annotators consistently surpasses 90%, substantially outperforming the level of agreement observed among human annotators.
@@
Alpaca-GPT-4-ZH (Peng et al., 2023) 4.16 4.87 4.51 3.99 4.75 4.37 4.83 4.68
COIG (Zhang et al., 2023a) 2.73 4.12 3.42 2.84 4.03 3.43 3.35 3.44
MOSS-SFT (Sun et al., 2024) 3.82 4.96 4.39 3.73 4.82 4.27 4.66 4.38
COIG-CQIA (Bai et al., 2024) 3.61 4.38 3.99 3.65 4.31 3.98 4.47 4.35
@@
Alpaca-GPT-4-ZH (Peng et al., 2023) 4.90 5.31 5.10 4.53 5.14 4.83 5.02 4.89
COIG (Zhang et al., 2023a) 2.82 4.41 3.62 3.11 4.38 3.75 3.48 3.66
MOSS-SFT (Sun et al., 2024) 4.71 5.26 4.98 4.55 5.17 4.86 5.06 4.90
COIG-CQIA (Bai et al., 2024) 4.72 5.13 4.92 4.51 5.02 4.76 4.73 4.51
@@
Alpaca-GPT-4-ZH (Peng et al., 2023) 2.77 4.28 3.53 2.64 4.12 3.38 4.21 4.28
COIG (Zhang et al., 2023a) 1.79 2.98 2.38 1.88 3.04 2.46 2.68 2.74
MOSS-SFT (Sun et al., 2024) 2.84 4.41 3.63 2.69 4.32 3.50 4.21 4.06
COIG-CQIA (Bai et al., 2024) 2.54 3.57 3.06 2.39 3.56 2.97 3.32 3.34
@@
Alpaca-GPT-4-ZH (Peng et al., 2023) 3.17 4.22 3.70 3.19 4.17 3.68 4.39 4.09
COIG (Zhang et al., 2023a) 1.70 2.81 2.25 1.95 2.80 2.38 2.69 2.65
MOSS-SFT (Sun et al., 2024) 2.84 4.21 3.53 2.73 4.19 3.46 4.14 4.23
COIG-CQIA (Bai et al., 2024) 2.69 3.58 3.14 2.64 3.56 3.10 3.36 3.13
@@
Alpaca-GPT-4-ZH (Peng et al., 2023) 5.64 5.73 5.68 5.34 5.73 5.54 5.44 5.36
COIG (Zhang et al., 2023a) 3.30 4.84 4.07 3.38 4.86 4.12 3.63 3.56
MOSS-SFT (Sun et al., 2024) 5.25 5.59 5.42 4.96 5.48 5.22 5.26 5.04
COIG-CQIA (Bai et al., 2024) 5.14 5.27 5.20 4.94 5.26 5.10 4.92 4.78
@@
Table 4: Performance of five LLMs on AlignBench and MT-Bench-zh after supervised fine-tuning on different datasets. “GPT-4o” and “DeepSeek-V3” indicate that responses were evaluated using GPT-4o and DeepSeek-V3.
When computing pairwise preference annotation agreement, we observe a small number of exceptional cases in which LLMs fail to strictly follow the annotation instructions and therefore do not produce a valid preference score, which prevents extraction of the corresponding preference label. We exclude such samples from the agreement computation. Importantly, these cases are extremely rare: among the 300 randomly selected samples, at most seven LLM-generated annotations fail to yield a valid preference label. Consequently, excluding these samples has a negligible impact on the overall results and does not affect the validity of our experimental conclusions.
6 Experiments
To assess TaP’s effectiveness and the quality of the preference datasets it constructs, we conducted supervised fine-tuning and preference fine-tuning (PPO/DPO) on LLMs using TaP-generated data.
6.1 Supervised Fine-tuning
To conduct supervised fine-tuning using the preference data generated by TaP, we utilized the prompts from the dataset along with GPT-4-generated responses. To ensure a balanced distribution of question types, we randomly selected one prompt and its corresponding response for each question type with multiple prompts. The resulting dataset comprises 19,840 samples generated by GPT-4 and 24,839 samples produced by DeepSeek-V2. A comparative summary of this supervised fine-tuning dataset, alongside other open-source Chinese datasets, is provided in Table 3.
6.2 Preference Fine-tuning
We adopt two dominant approaches for preference fine-tuning using the dataset constructed by TaP: PPO and DPO. For comparison with other open-source preference datasets, we utilize Huozi-RLHF (Huozi-Team, 2024) and Chinese-DPO-Pairs [^9] https://huggingface.co/datasets/wenbopan/Chinese-dpo-pairs for preference fine-tuning. Further details on the PPO and DPO setups are provided in Appendix D.1.
@@ 6.2 Preference Fine-tuning
Qwen2.5-3B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 4.02 5.23 4.63 3.88 5.04 4.46 5.03 4.94
@@
Qwen2.5-3B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 5.15 5.82 5.49 4.65 5.50 5.08 5.72 5.24
@@
Qwen2.5-7B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 4.53 5.39 4.96 4.37 5.34 4.85 5.25 4.98
@@
Qwen2.5-7B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 6.24 6.34 6.29 5.83 6.00 5.92 6.10 5.85
@@
Llama-3.1-8B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 3.52 5.19 4.36 3.29 5.06 4.17 4.85 4.58
@@
Llama-3.1-8B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 3.70 5.22 4.46 3.33 4.86 4.09 5.19 4.84
@@
Gemma-2-9B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 3.93 5.22 4.58 3.54 5.05 4.29 4.77 4.62
@@
Gemma-2-9B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 4.18 5.23 4.70 3.77 4.93 4.35 5.39 5.01
@@
Qwen2.5-14B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 5.75 5.90 5.82 5.64 5.81 5.73 5.41 5.48
@@
Qwen2.5-14B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 6.94 6.64 6.79 6.41 6.28 6.35 6.25 6.11
@@
Qwen2.5-3B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 4.13 4.88 4.51 4.44 5.36 4.90 5.25 5.20
@@
Qwen2.5-3B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 4.61 5.27 4.94 4.62 5.55 5.08 5.86 5.76
@@
Qwen2.5-7B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 4.71 5.56 5.13 4.46 5.43 4.95 5.21 5.19
@@
Qwen2.5-7B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 6.24 6.44 6.34 5.67 5.98 5.82 6.25 6.19
@@
Llama-3.1-8B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 3.72 5.27 4.50 3.41 4.97 4.19 4.93 4.89
@@
Llama-3.1-8B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 3.94 5.25 4.60 3.61 4.86 4.24 5.25 5.04
@@
Gemma-2-9B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 3.75 4.75 4.25 3.81 5.14 4.47 4.87 4.85
@@
Gemma-2-9B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 3.85 5.20 4.53 3.56 4.96 4.26 5.27 5.11
@@
Qwen2.5-14B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 5.61 5.90 5.75 5.69 5.93 5.81 5.78 5.78
@@
Qwen2.5-14B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 6.61 6.50 6.55 6.42 6.34 6.38 6.41 6.51